### v1

@@ Abstract
Affective polarization and its accompanying cleavage-based sorting drives incivility and contentiousness around climate change and other science-related issues. Looking at the COVID-19 period, we study cross-domain spillover of incivility and contentiousness in public engagements with climate change and climate science on Twitter and Reddit. We find strong evidence of the signatures of affective polarization surrounding COVID-19 spilling into the climate change domain. Across different social media systems, COVID-19 content is associated with incivility and contentiousness in climate discussions. These patterns of increased antagonism were responsive to pandemic events that made the link between science and public policy more salient. We also show that the observed spillover activated along pre-pandemic political cleavages, specifically anti-internationalist populist beliefs, that linked climate policy opposition to vaccine hesitancy. Our findings highlight the dangers of entrenched cross-domain polarization manifesting as spillover of antagonistic behavior.
@@ Introduction
There is a virtual consensus in the scientific community that climate change is caused by human activity (Myers et al. 2021; Cook et al. 2013), and an overwhelming majority of scientists across disciplines agree that fundamental changes are needed to address the climate crisis (Dablander et al. 2024). Despite this, a non-trivial minority of the public remains firmly skeptical of climate science and opposed to stronger climate policies (Leiserowitz et al. 2023). Notably, they are often vocal about their skepticism and opposition on social media platforms like Twitter (Falkenberg et al. 2022).
@@
Evidence-based debates over findings are the cornerstone of the scientific endeavor, but unsubstantiated claims that run counter to the scientific consensus and questions about the integrity of scientists are linked to the delegitimization of science and lower trust for scientists among the public (Leiserowitz et al. 2013), as is the presence of disagreements and incivility in media coverage about science (Chinn and Hart 2022). Just as important, research into the public’s antagonistic engagement with science-related issues finds extensive evidence of a host of additional deleterious phenomena that have been associated with increased social and political polarization, including motivated reasoning (Druckman and McGrath 2019), partisan sorting (Chen et al. 2021), and the politicization of science (Druckman 2017).
@@
In this study, we explore public engagement with climate change and climate science on social media, focusing on how manifestations of affective polarization over the climate issue – animosity toward members of the political outgroup (Iyengar and Krupenkin 2018) – are affected by spillover from politically divisive events in other domains of science. We examine patterns of incivility and contentiousness, which are harmful outcomes of affective polarization (Mason 2015). Incivility, which ranges from incensed discussions to outright rude critiques and name-calling (Anderson et al. 2014), intensifies polarization (Brundidge and Garrett 2024; Chen, Song and Guo 2024). It is associated with contributing to the dehumanization of outgroup members (Moore-Berg et al. 2020) and the overall decrease in the quality of the information environment by, e.g., reducing individuals’ trust (Mutz and Reeves 2005) and willingness to engage (Ng et al. 2020). Public contentiousness over scientific findings is further problematic in the context of climate change as it runs counter to the scientific community’s overwhelming agreement regarding the anthropogenic sources of climate change.
@@
More generally, spillover across different societal issues is itself evidence of polarization, which has implications for science communication and policymaking, as it indicates social or political cleavages are becoming more entrenched (Baldassarri and Gelman 2008; Kozlowski and Murphy 2021). To study spillover, which is both a mechanism and outcome of polarization, we look at the COVID-19 period, during which the severity of the pandemic increased the visibility and importance of communication between the general public and the scientific community. This period saw high levels of politicization and polarization toward COVID-19 (Green et al. 2020; Hart, Chinn and Soroka 2020), which likely contributed to the diverging trust in science among different segments of the population (Radrizzani et al. 2023; Kerr, Panagopoulos and Van Der Linden 2021). The impact of the COVID-19 shock therefore provides us an opportunity to examine important patterns in how the public engages with science across different issues, as defining and novel events from the pandemic can reveal evidence of spillover from COVID-19 to climate science discussions.
We explore the following questions about how cross-domain spillover leads to greater polarization in how the public engages with science. How do large scale shocks to the role of science in society change how the public engages with scientific issues across different domains? Do negative aspects of COVID-19 science spill over to public engagement with climate science? What are the conceptual pathways through which these spillovers travel?
@@
There is a growing tendency for issues to become linked through their association with partisan positions, whereby individuals’ partisanship predict their attitudes toward issues, even to the extent that their positions toward different issues become predictive of one another (Kozlowski and Murphy 2021). With the constant spotlight on the role of science in policymaking during the pandemic, we expect public tendencies surrounding COVID-19 to spill into how the public engages with climate science. Specifically, given the salience of polarized elite political cues surrounding COVID-19 (Green et al. 2020), we expect increased incivility in climate discussions that also involve COVID-19 topics. We also expect the combination of diverging trust toward the institution of science and the increasing inclination for individuals to do their “own research” (Chinn and Hasell 2023) to lead to greater contentiousness in public debates about climate science issues.
@@
We approach our research questions by studying antagonistic public engagement with climate change-related content on social media, including those that explicitly engage with published climate science literature. We focus on Twitter, which, at least until 2022, was an important space for public contestations over social and political issues (Theocharis 2015). Moreover, within the scientific community, there had been a generally shared notion that Twitter is the primary social media platform for high salience science communication (Hammer, Boender and Thomas 2021), which makes it an interesting case for questions about the science-public interface (Walter, Lörcher and Brüggemann 2019). We also examine Reddit to supplement our findings. We show that the presence of COVID-19 topics is strongly predictive of climate change conversations taking a turn toward increased incivility and contentiousness, that these patterns respond to landmark events in the COVID-19 domain, and that incivility spillover tends to activate along existing sources of political polarization.
@@ Results
We study the impact of cross-domain spillover between public science issues in two main ways. First, we consider how concepts from one domain are used in another by looking at patterns of incivility and contentiousness when COVID-19 enters the climate change narrative. By examining this intersection, we are able to uncover shifts in polarization that arise when global crises intertwine with wider scientific debates. Second, we look at how antagonistic communication patterns in climate change and climate science align with important COVID-19 events over time.
@@
To assess whether there is spillover between the two public science domains of COVID-19 and climate change, we compared the incivility of all pandemic period (March 20, 2020 onward) climate posts based on whether they contained COVID-19 content. We also looked at climate posts that specifically invoked Anthony Fauci, who, as then-Chief Medical Advisor to the President of the U.S., polarized COVID-19 discussions by being emblematic of how the institution of science impacts politics (Chen et al. 2023). To measure incivility, we used toxicity detection from Perspective API (Jigsaw 2017), which provides the probability a post is “a rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.” This operationalization maps directly onto the disengaging effects of incivility that we want to capture. From examples of tweets identified to be toxic (Table 1), we see that both COVID-19 and Fauci as topics capture polarized, vitriolic debates on the science involved.
$P$(toxic) Tweet Text
0.944 Are people really that stupid? Do they really believe this crap? Climate change is a bunch of bullshit so is Dr. Fauci. It’s just pathetic. And this figures just every now and then they’re gonna make people lock down and do the mask bullshit again. Weak minded fucking sheep. [pouting-face emoji]
@@
0.654 When you find out climate change scientists are all like Dr. Fauci. Phoney flakes who don’t know shit from shinola.
@@
Table 1: Examples of uncivil climate tweets that reference Anthony Fauci. Tweet texts have been edited for typos and formatting.
Across all three data sets – general climate Twitter, climate science Twitter, and climate Reddit – we find evidence of incivility spillover between climate change and COVID-19. Estimates from linear probability models of toxicity in climate posts show that posts invoking COVID-19 and Fauci have a significantly higher probability of being toxic, with the exception of climate science tweets referencing only COVID-19 (Figure 1.A). [^1] Our results are robust to estimates from fractional logistic regression models (Papke and Wooldridge 2008). Among general climate tweets, these results hold when account fixed effects are included, meaning the observed pattern of COVID-19 posts being more toxic is not simply due to different individuals choosing different things to talk about, but is attributable to an individual-level behavioral tendency to be more toxic when discussing COVID-19.222We do not fit account fixed effects models for the scientific tweets data because there is not enough within-user variation in our predictors, nor for the Reddit data because it does not contain user information.
Our results are robust to controlling for whether the post mentions international organizations or figures who are targets of conspiracy theories (a point we return to in Section 2.3); whether the post is the first post of a conversation, which has a higher chance of being a news or scientific article posted by an institutional account; and whether the posting account was active before the COVID-19 period because we expect Twitter to have attracted new users during the pandemic who are more uncivil than existing ones.
@@
Figure 1: Presence of COVID-19 content and incivility of climate posts. Panel A shows coefficient estimates from linear probability model of toxicity in climate posts across three social media systems. Panel B shows coefficient estimates from Cox proportional hazards models of toxicity onset in climate conversations on Twitter. Panel C shows temporal trends in daily climate tweet toxicity by difference in content. Panel C’s inset shows the temporal trends for climate science tweets and Reddit comments.
Next, we leverage the sequential structure of conversations to show that general climate tweets that include references to COVID-19 tend to incite greater toxicity in their replies. From Cox proportional hazards models of toxicity onset in general climate and climate science conversations, we find that the number of replies until one exceeds 0.5 toxicity probability is significantly lower in general climate conversation that begin with a tweet mentioning COVID-19 (Figure 1.B). At any given time, replies to these conversations are 9% more likely to devolve into incivility compared to conversations that do not begin with COVID-19 content. This pattern however does not hold for climate science conversations. Further, corroborating results from our post-level models, we find that toxicity onset in both general climate and climate science conversations are likely to be concurrent to mentions of COVID-19 or Fauci.
@@
Finally, we find that the tendency for users to invoke COVID-19 in more toxic posts aligns with important pandemic events (Figure 1.C). Here, we estimated toxicity trends for climate posts with and without COVID-19 content in all three data sets using the BEAST time series method (Zhao et al. 2019), a trend and change point estimation method designed for noisy and cyclical data, such as social media patterns (Reuning, Whitesell and Hannah 2022). First, as in the linear models, the presence of COVID-19 content is strongly predictive of toxicity in general climate discussions on both Twitter and Reddit, but not in posts that explicitly link to climate science. Second, increases in the tendency for invoking COVID-19 in toxic climate discussions clearly correspond to times when the link between COVID-19 science and public policy became more salient, such as the lead up to the 2020 U.S. presidential elections and the onset of the Delta variant and its associated policy responses in early summer 2021. Third, COVID-19 climate tweets initially start at approximately 0.1 toxicity probability, which is the same as general climate tweets, but quickly spike within the first three months. This temporal pattern is indicative of partisan sorting, whereby new issues enter the public consciousness as nonpartisan but quickly become divisive and politically charged as a result of elite-led polarization, which was strong during the COVID-19 period (Green et al. 2020).
@@
Figure 2: Presence of COVID-19 content and probability of disagreement between climate conversation participants. Panel A shows coefficient estimates from linear probability models of disagreement in different types of climate conversations. Panel B shows temporal trends in weekly average climate conversation disagreement probability by whether the first tweet mentions COVID-19.
@@
On the whole, we find that climate conversations that engage with COVID-19 topics are significantly more contentious than those that do not. First, we fit linear probability models of response disagreement in all message-pairs to see whether linking climate change to COVID-19 resulted in higher levels of disagreement and whether individuals invoked COVID-19 to disagree about climate change (Figure 2.A). We find evidence for both in conversations about climate science publications, as COVID-19 content in root and response tweets were both associated with higher levels of response disagreement. Table 2 shows examples of both types of contentiousness spillover. In general climate conversations, we only find evidence for the use of COVID-19 content to disagree with others.
For our BEAST time series analysis (Figure 2.B), which focuses on whether the conversation is related to COVID-19 (measured by whether the root tweet contains COVID-19 keywords), we calculated conversation-level disagreement by averaging over all message-pairs in the conversation. We first find that disagreement with climate change posts immediately dropped with COVID-19 onset – a pattern also observed with the incivility of climate posts – which indirectly substantiates the “finite pool of attention” hypothesis that the pandemic drew the public’s attention away from the climate issue (Sisco et al. 2023). At the same time, however, climate conversations that relate to COVID-19 grew sharply in disagreement levels, and by the end of our observation period in August 2021, the level of disagreement in climate conversations had rebounded to above pre-pandemic highs. This pattern was even stronger in conversations directly engaging with climate science publications, with COVID-19 topics immediately exceeding pre-pandemic levels of disagreement.
$P$(disagree) Tweet-Reply Pairs
@@
0.801 Excited collaboration with 43 authors from 27 institutes shows COVID-19 caused the largest decrease of CO2 emissions since 1900. Larger than that of WWII.
@@
0.918 China’s iron & steel industry is a major source of air pollution. In 2015 it brought in tighter standards. From 2014 to 2018 (only 4 years!): Particulate matter (PM) fell by 47%. SO2 fell by 42%. Change can happen quickly.
@@
Table 2: Examples of high disagreement responses in climate science conversations that reference COVID-19. Tweet texts have been edited for typos and formatting.

### v2

@@ Abstract
Affective polarization and political sorting drive public antagonism around issues at the science-policy nexus. Looking at the COVID-19 period, we study cross-domain spillover of incivility and contentiousness in public engagements with climate change and public health on Twitter and Reddit. We find strong evidence of the signatures of affective polarization surrounding COVID-19 spilling into the climate change domain. Across different social media systems, COVID-19 content is associated with incivility and contentiousness in climate discussions. These patterns of increased antagonism were responsive to pandemic events that made the link between science and public policy more salient. The observed spillover activated along pre-pandemic political cleavages, specifically anti-internationalist populist beliefs, that linked climate policy opposition to vaccine hesitancy. Our findings show how affective polarization in public engagement with science becomes entrenched across science policy domains.
@@ Introduction
There is a virtual consensus in the scientific community that climate change is caused by human activity (Myers et al. 2021; Cook et al. 2013), and an overwhelming majority of scientists across disciplines agree that fundamental changes are needed to address the climate crisis (Dablander et al. 2024). Despite this, a non-trivial minority of the public firmly remains skeptical of climate science and opposed to stronger climate policies (Leiserowitz et al. 2023). Notably, they are vocal about their skepticism and opposition toward climate science on social media platforms like Twitter (Falkenberg et al. 2022; Rojas et al. 2024), often in a highly antagonistic manner (Xia, Chen and Kivelä 2021). Similar phenomena exist in the public health realm, which became widespread during the COVID-19 pandemic when antagonistic anti-vaccine and anti-science sentiments were rampant on social media (Hsieh 2025).
@@
Evidence-based debates over findings are the cornerstone of the scientific endeavor, but unsubstantiated claims that run counter to the scientific consensus and questions about the integrity of scientists are linked to the delegitimization of science and lower trust for scientists among the public (Leiserowitz et al. 2013), as is the presence of disagreements and incivility in media coverage about science (Chinn and Hart 2022). Just as important, research into the public’s contemporary engagement with issues at the science-policy nexus finds a host of deleterious phenomena associated with social and political polarization, including motivated reasoning (Druckman and McGrath 2019), partisan sorting (Chen et al. 2021), and the politicization of science (Druckman 2017), all of which have severe implications for science communication and policymaking.
@@
In this study, we explore the public’s antagonistic engagement with climate science and public health, two critical issues at the science-policy nexus. We focus on how manifestations of affective polarization – animosity toward members of the political outgroup (Iyengar and Krupenkin 2018) – spill across different issue domains. Specifically, we examine patterns of incivility and contentiousness. Incivility, which ranges from incensed discussions to outright rude critiques and name-calling (Anderson et al. 2014), intensifies polarization (Brundidge and Garrett 2024; Chen, Song and Guo 2024). It is associated with contributing to the dehumanization of outgroup members (Moore-Berg et al. 2020), and the overall decrease in the quality of the information environment by, e.g., reducing individuals’ trust (Mutz and Reeves 2005) and willingness to engage (Ng et al. 2020). Public contentiousness over scientific findings in these areas is further problematic as it runs counter to the scientific community’s overwhelming agreement regarding the anthropogenic causes of climate change and the effectiveness of vaccines for public health.
We conceptualize spillover of antagonism between societal issues as a manifestation of affective polarization resulting from individuals sorting into entrenched political groups (Mason 2015). There is a growing tendency for issues to become linked through their association with partisan positions, whereby individuals’ attitudes toward different issues become predictive of each other (Kozlowski and Murphy 2021). As more issues become aligned in this manner, entrenched group-based cleavages become salient cues for belief formation and political behavior (Baldassarri and Gelman 2008), including how one understands new policy issues and engages with their outgroup. When a highly aligned system is also affectively polarized, antagonistic communication cues and outgroup mistrust in one domain should easily travel to another. Following this mechanism, linkages between issues that share overlapping sources of political cleavage can also become a pathway which political extremism, e.g., politically-linked conspiracy beliefs, travels through (Miller 2020; Enders et al. 2021), a particularly salient concern in the public health domain (Erokhin, Yosipof and Komendantova 2022).
@@
To study spillover between climate change and public health, we look at the COVID-19 period. During this period, the severity of the pandemic increased the visibility and importance of communication between the general public and the scientific community. This period saw high levels of politicization and polarization toward COVID-19 (Green et al. 2020; Hart, Chinn and Soroka 2020), which likely contributed to the diverging trust in science among different segments of the population (Radrizzani et al. 2023; Kerr, Panagopoulos and Van Der Linden 2021). The impact of the COVID-19 shock therefore provides us an opportunity to examine important patterns in how the public engages with science across different issues, as novel events from the pandemic can reveal evidence of spillover from public health to climate science discussions.
@@
We explore the following questions about how cross-domain spillover leads to greater polarization in how the public engages with science. How do large scale shocks to the role of science in society change how the public engages with scientific issues across different domains? Do negative aspects of COVID-19 science spill over to public engagement with climate science? What are the conceptual pathways through which these spillovers travel? With the constant spotlight on the role of science in policymaking during the pandemic, we expect public tendencies surrounding COVID-19 to spill into how the public engages with climate science. Specifically, given the salience of polarized elite political cues surrounding COVID-19 (Green et al. 2020), we expect increased incivility in climate discussions that also involve COVID-19 topics. We also expect the combination of diverging trust toward the institution of science and the increasing inclination for individuals to do their “own research” (Chinn and Hasell 2023) to lead to greater contentiousness in public debates about climate science issues.
@@
We approach our research questions by studying antagonistic public engagement with COVID-19 and climate change-related content on social media, including those that explicitly engage with published climate science literature. We focus on Twitter, which, at least until 2022, was an important space for public contestations over social and political issues (Theocharis 2015). Moreover, within the scientific community, there had been a generally shared notion that Twitter is the primary social media platform for high salience science communication (Hammer, Boender and Thomas 2021), which makes it an important case for questions about the science-public interface (Walter, Lörcher and Brüggemann 2019). We also examine Reddit discussions to supplement our findings. We show that the presence of COVID-19 content is strongly predictive of climate change conversations taking a turn toward increased incivility and contentiousness, that these patterns respond to landmark events in the COVID-19 domain, and that incivility spillover tends to activate along existing sources of political polarization.
@@ Results
We study the impact of cross-domain spillover between science policy issues by examining how patterns of incivility and contentiousness align with COVID-19 entering the climate change narrative. First, we look at how these communication patterns vary with the onset of important COVID-19 events over time. Second, by fitting statistical models for co-occurrence and onset, we explore in further detail the conditions under which COVID-19 impacts climate discussions. By examining this intersection, we are able to uncover shifts in polarization that arise when global crises intertwine with wider scientific debates.
@@
Tweet Text
Are people really that stupid? Do they really believe this crap? Climate change is a bunch of bullshit so is Dr. Fauci. It’s just pathetic. And this figures just every now and then they’re gonna make people lock down and do the mask bullshit again. Weak minded fucking sheep.
@@
When you find out climate change scientists are all like Dr. Fauci. Phoney flakes who don’t know shit from shinola.
@@
Table 1: Examples of uncivil climate tweets that reference Anthony Fauci. Toxicity probabilities of these tweets exceed 0.65. Tweet texts have been edited for typos and formatting.
To assess whether there is spillover between the two science policy domains of COVID-19 and climate change, we compared the incivility of all pandemic period (March 20, 2020 onward) climate posts based on whether they contained COVID-19 content. We also looked at climate posts that specifically invoked Anthony Fauci, who, as then-Chief Medical Advisor to the President of the U.S., polarized COVID-19 discussions by being emblematic of how the institution of science impacts politics (Chen et al. 2023). To measure incivility, we used toxicity detection from Perspective API (Jigsaw 2017), which provides the probability a post is “a rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.” This operationalization maps directly onto the disengaging effects of incivility that we want to capture. From examples of toxic tweets (Table 1), we see that both COVID-19 and Fauci content were used in polarized, vitriolic debates on climate science.
@@
Figure 1: Presence of COVID-19 content and incivility of climate posts. A. Temporal trends in daily climate tweet toxicity by difference in content. Inset shows the temporal trends for climate science tweets and Reddit comments. B. Coefficient estimates from linear probability model of toxicity in climate posts across three social media systems. C. Coefficient estimates from Cox proportional hazards models of toxicity onset in climate conversations on Twitter.
@@
First looking at broader temporal trends, we find a generally strong tendency for users to invoke COVID-19 in more toxic posts, a pattern which is aligned with important pandemic events (Fig. 1.A). Here, we estimated toxicity trends for climate posts with and without COVID-19 and Fauci content in all three data sets – general climate Twitter, climate science Twitter, and climate Reddit – using the Bayesian Estimator of Abrupt change, Seasonal change, and Trend (BEAST) time series method (Zhao et al. 2019), a trend and change point estimation model designed for noisy and cyclical data, such as social media patterns (Reuning, Whitesell and Hannah 2022).
We find three notable patterns. First, the presence of COVID-19 content is strongly predictive of toxicity in climate discussions on both Twitter and Reddit, with the exception of posts that explicitly link to climate science. This pattern is even stronger for Fauci content, which we only examined in general climate Twitter discussions due to their low presence in the other data sets. Second, increases in the tendency for invoking COVID-19 in toxic climate discussions clearly correspond to times when the link between COVID-19 science and public policy became more salient, such as the lead up to the 2020 U.S. presidential elections and the onset of the Delta variant and its associated policy responses in early summer 2021. Third, COVID-19 climate tweets initially start at approximately 0.1 toxicity probability (March–May), which is the same as general climate tweets that do not contain COVID-19 content, but quickly spike and diverge from the baseline within the first three months (June onward). This temporal pattern is indicative of partisan sorting, whereby new issues enter the public consciousness as nonpartisan ones, but quickly become divisive and politically charged as a result of elite-led polarization (Hetherington 2001), which was strong during this period (Green et al. 2020).
Next, we look more closely at content toxicity patterns after the initial incivility escalation period (Fig. 1.B). Across all three data sets, we find consistent evidence of incivility spillover between climate change and COVID-19 during September 2020–August 2021. Estimates from linear probability models of toxicity in climate posts show that posts invoking COVID-19 and Fauci have a significantly higher probability of being toxic, although climate science tweets very rarely contained Fauci content. Among general climate tweets, these results hold when account fixed effects are included, meaning the observed pattern of COVID-19 posts being more toxic is not simply due to different individuals choosing different things to talk about, but is attributable to an individual-level behavioral tendency to be more toxic when discussing COVID-19. We do not fit account fixed effects models for the scientific tweets data because there is not enough within-user variation in our predictors, nor for the Reddit data because it does not contain user information.
Our results are robust to controlling for whether the post mentions international organizations or public figures who are targets of conspiracy theories, and whether the post contains obstructionist or contrarian information about climate change (two discussions we return to later); whether the post is the first post of a conversation, which has a higher chance of being a news or scientific article posted by an institutional account; and whether the posting account was active before the COVID-19 period because we expect Twitter to have attracted new users during the pandemic who are more uncivil than existing ones. Our results also hold when we use fractional logistic regression models (Papke and Wooldridge 2008), presented in Fig. S6.1.
Finally, we leverage the sequential structure of conversations to show that general climate tweets that include references to COVID-19 tend to incite greater toxicity in their replies. From Cox proportional hazards models of toxicity onset in general climate and climate science conversations, we find that the number of replies until one exceeds 0.5 toxicity probability is significantly lower in general climate conversation that begin with a tweet mentioning COVID-19 (Fig. 1.C). At any given time, replies to these conversations have approximately 14% higher odds to devolve into incivility compared to conversations that do not begin with COVID-19 content. This pattern however does not hold for climate science conversations. Further, corroborating results from our post-level models, we find that toxicity onset in both general climate and climate science conversations are likely to be concurrent to mentions of COVID-19 or Fauci.
@@
For our BEAST time series analysis (Fig. 2.A), we calculated conversation-level disagreement by averaging over all tweet-reply pairs in the conversation, fitting separate models by whether the root tweet of the conversation contains COVID-19 keywords. We first find that disagreement with climate change posts immediately dropped with COVID-19 onset – a pattern also observed with the incivility of climate posts – which indirectly substantiates the “finite pool of attention” hypothesis that the pandemic drew the public’s attention away from the climate issue (Sisco et al. 2023). At the same time, however, climate conversations that relate to COVID-19 grew sharply in disagreement levels, exhibiting a spike similar to content toxicity starting from September 2020. By the end of our observation period in August 2021, the level of disagreement in climate conversations had rebounded to above pre-pandemic highs. This pattern was even stronger in conversations directly engaging with climate science publications, with COVID-19 topics immediately exceeding pre-pandemic levels of disagreement.
Figure 2: Presence of COVID-19 content and probability of disagreement between climate conversation participants. A. Temporal trends in weekly average climate conversation disagreement probability by whether the first tweet mentions COVID-19. B. Coefficient estimates from linear probability models of disagreement in different types of climate conversations. C. Joint distribution of comments having at least 0.8 contrarian probability and their responses having at least 0.5 disagreement probability (left). Increase in probability of a response containing obstructionist content when disagreeing, versus not disagreeing, with different types of comments (right).
Next, we fit linear probability models with all tweet-reply pairs during the September 2020–August 2021 period to better understand why COVID-19 conversation are more contentious. Specifically we examine whether linking climate change to COVID-19 resulted in higher levels of disagreement or whether individuals invoked COVID-19 to disagree about climate change, both of which we show examples of in Table 2. We find evidence of both types of contentiousness spillover in conversations about climate science publications, as COVID-19 content in both the initial tweet and its response were associated with higher levels of response disagreement (Fig. 2.B). On the other hand, in general climate conversations, we only find evidence for the use of COVID-19 content to disagree with others.
Tweet-Reply Pairs
@@
Excited collaboration with 43 authors from 27 institutes shows COVID-19 caused the largest decrease of CO2 emissions since 1900. Larger than that of WWII.
@@
China’s iron & steel industry is a major source of air pollution. In 2015 it brought in tighter standards. From 2014 to 2018 (only 4 years!): Particulate matter (PM) fell by 47%. SO2 fell by 42%. Change can happen quickly.
@@
Table 2: Examples of high disagreement responses in climate science conversations that reference COVID-19. Disagreement probabilities of these tweet-pairs exceed 0.80. Tweet texts have been edited for typos and formatting.
2.2.1 Disagreement and Climate Obstructionist Content
Given these results, we next consider the relationship between disagreement and harmful interactions, as disagreement in itself can also be a sign of healthy discourse – indeed we find that tweets containing climate obstructionist content are approximately 10% more likely to be disagreed with. Here, we proceed with the notion that because there is overwhelming agreement among scientists in both climate science (Myers et al. 2021) and climate policy science (Dablander et al. 2024), contrarian positions should be understood as a form of harmful climate obstruction (Rojas et al. 2024). We explore this by examining the extent to which disagreement in tweet-reply pairs varies by the presence of contrarian claims, estimated at the individual tweet level using the Augmented Computer Assisted Recognition of Denial and Skepticism (CARDS) model which has been specifically tuned for this task (Rojas et al. 2024).
First, we plot the joint distribution of whether a tweet contains contrarian claims and whether its response disagrees with it (Fig. 2.C, left). This shows that approximately 90% of responses – in both general climate and climate science data sets – are directed at non-obstructionist content. Of these, approximately 30% are disagreements. We then compared the obstructionist content of responses based on whether they disagree with non-obstructionist tweets (Fig. 2.C, right). In both data sets, we find that there is a 10–15% higher chance that disagreements with non-obstructionist content invoke contrarian claims, which is a significantly higher rate than all other combinations of tweet-reply pairs.
These results together show that while disagreement is not in itself harmful, in the context of climate science and policy discussions where there is overwhelming scientific agreement, disagreement tends to draw on unsubstantiated contrarian challenges, previously demonstrated to have harmful impacts on public trust toward science (Leiserowitz et al. 2013).